# 2cm2cm1.5cm1.5cm

### v1

@@ Appendix B Majorization inequality for the volume bound
The functional $W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})=\prod\limits_{i\in[m]}x_{i}^{\delta_{i}x_{i}^{2}}$ under the constraints
  • $\frac{1}{d^{\prime}+1}\leq x_{i}\leq 1\quad\forall\,i\in[m],$
@@
$\ln W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{d+1}\ln(d+1).$
Note that the function $\ln W$ is continuous on the set in $\mathbb{R}^{2m}$ defined by our constraints. Hence, it attains its maximum. For every $x=(x_{1},\ldots,x_{m})$ with $\frac{1}{d+1}\leq x_{i}\leq 1$ for all $i\in[m]$, let $\ln W_{x}(\delta_{1},\dots,\delta_{m})$ be the function
@@
If $\delta_{\lambda}^{*}$ is the decreasing rearrangement of $\delta_{\lambda}$, we can assume without loss of generality that $\delta_{\lambda}=\delta_{\lambda}^{*}$. Set
@@
$\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}(x)=\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}=\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{d+1}\ln(d+1).$
@@ Acknowledgements
We thank Alexander Litvak for the fruitful conversations. We are also grateful to Susanna Dann, who organized the ’Analysis and Convex Geometry Week at UniAndes’ in Bogotá in 2025, where the initial steps of this work were made.
G. I. is supported by Projeto Paz and by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior) - Brasil, grant number 23038.015548/2016-06. Z.L. and M.N. is partially supported by the NRDI research grant K147544. M.N. is also supported by the NRDI research grant K131529 as well as the ELTE TKP 2021-NKTA-62 funding scheme. Á. S. is funded by the grant NKFIH 150151, a project that has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.

### v2

@@ Appendix B Majorization inequality for the volume bound
The following lemma yields (5.3) with the substitution $\left|P_{F}u_{i}\right|^{2}=x_{i}$ and $d=d^{\prime}.$ We leave it to the reader that the constraints on $x_{i}$ are satisfied.
@@
The functional $W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})=\prod\limits_{i\in[m]}x_{i}^{\frac{\delta_{i}x_{i}}{2}}$ under the constraints
  • $\frac{1}{d+1}\leq x_{i}\leq 1\quad\forall\,i\in[m],$
@@
$\ln W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{2(d+1)}\ln(d+1).$
Note that the function $\ln W$ is continuous on the compact set in $\mathbb{R}^{2m}$ defined by our constraints. Hence, it attains its maximum. For every $x=(x_{1},\ldots,x_{m})$ with $\frac{1}{d+1}\leq x_{i}\leq 1$ for all $i\in[m]$, let $\ln W_{x}(\delta_{1},\dots,\delta_{m})$ be the function
@@
Set
@@
$\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}(x)=\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}=\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{2(d+1)}\ln(d+1).$
@@ Acknowledgements
We thank Alexander Litvak and Roman Karasev for the fruitful conversations. We are also grateful to Susanna Dann, who organized the ‘Analysis and Convex Geometry Week at UniAndes’ in Bogotá in 2025, where the initial steps of this work were made.
G. I. is supported by Projeto Paz and by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior) - Brasil, grant number 23038.015548/2016-06. Z. L. and M. N. are partially supported by the NRDI research grant K147544. M. N. is also supported by the NRDI research grant K131529 as well as the ELTE TKP 2021-NKTA-62 funding scheme. Z. L. is supported also by the ERC Advanced Grant “ERMiD”, Á. S. is funded by the grant NKFIH 150151, a project that has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.